OpenStudy (anonymous):

confirm that f and g are inverses by showing that f(g(x))=x and g(f(x))=x. f(x)=2/x and g(x)=2/x

OpenStudy (anonymous):

To give you two different examples, let f and g be functions defined as follows: f(x)=x + 5/6 g(y)=6y - 5 (I took the liberty of using y in the definition for g) Then you verify as follows: g(f(x)) = g(x + 5/6) = 6(x + 5/6)-5 = 6x + 5 - 5 = 6x In this case, f and g are not inverses of each other. Now, if you were to find that g(f(x)) = x, you are halfway home. The other half is showing that f(g(y)) = y. got it from yahoo answers